# REPS: Recycled Entropy Packet Spraying for Adaptive Load Balancing and Failure Mitigation

### v5

@@ Abstract.
submissionid: 52copyright: nonedoi: isbn: price:
@@ 1. Introduction
Modern AI training cluster networks draw on both high-performance computing (HPC) and cloud-native architectures (Hoefler et al., 2022b; Qian et al., 2024; Gangidi et al., 2024). These systems rely on Remote Direct Memory Access (RDMA) for low-latency, high-throughput communication. InfiniBand, a specialized RDMA transport, is widely used for its performance (269, 2004). More recently, operators have adopted commodity Ethernet fabrics, especially RoCEv2 (Association, 2024), which enable RDMA over standard hardware using off-the-shelf switches and NICs to reduce costs. However, as training clusters scale from around 10,000 endpoints (enough to train GPT-4 or Llama 3–scale models (Achiam et al., 2023; Dubey et al., 2024)) to over 100,000 nodes, both interconnects face fundamental limits in sustaining peak bandwidth due to:
@@
Thus, the community recognized the need for network stacks tailored to distributed training traffic while remaining compatible with commodity datacenter Ethernet infrastructure (Gangidi et al., 2024). Such proposals include SRD by Amazon (Shalev et al., 2020), Falcon by Google (goo, 2023), TTPoE by Tesla (Motors, 2024), and the upcoming Ultra Ethernet (UE) (Consortium, 2024), developed in collaboration between major tech players. Key open questions in these proposals are how to address load balancing and mitigate link failures.
@@
Our key insight is that the challenges of spraying-based load balancing can be addressed by adaptive packet spraying, paired with a transport layer that natively supports out-of-order packet delivery, as demonstrated in SDR (Shalev et al., 2020), UE (Lu et al., 2018), Falcon (goo, 2023) and recent research work (Dixit et al., 2013; NVIDIA, 2024; Gangidi et al., 2024). Based on this insight, we design and contribute to the Ultra Ethernet Consortium (UEC) our decentralized load-balancing scheme, Recycled Entropy Packet Spraying aka REPS. REPS caches ”good” network paths in a circular buffer and quickly recovers (within a few RTTs) from network failures by adaptively discovering or freezing network paths.
REPS does not require any specific hardware support from switches beyond ECMP-based header hashing and ECN, which are standard features in modern switches (Hoefler et al., 2023; Nvidia, 2024). REPS requires only $\approx 25$ bytes of state per-connection regardless of the topology size, whereas MPTCP requires 368 extra bytes for 8 sub-flows (Lu et al., 2018).
We extensively evaluate REPS in simulation and augment the simulation findings with real hardware results in Section 4. We deploy REPS in a cluster with modified FPGA-based RDMA-capable NICs. In large-scale network simulations, REPS consistently outperforms state-of-the-art load balancing algorithms. REPS outperforms ECMP and OPS by up to $6\times$ and $1.25\times$ in symmetric networks, and by up to $4.5\times$ and $2\times$ in asymmetric networks, outperforming OPS by as much as $100\times$ during short-term transient link failures.
@@ 2. REPS Building Blocks
In this section, we introduce the fundamental building blocks necessary to understand REPS logic. We first describe key congestion signals, such as ECN and packet loss. We then introduce essential terms and concepts related to load balancing, including ECMP and EV. These concepts are largely established, but we revisit them here to provide clear terminology and a basis for the discussion that follows.
@@ 2.1. Congestion Signals
has long been a key indicator of severe congestion in networks (Floyd and Jacobson, 1993; Nichols and Jacobson, 2012). However, using packet losses as the sole signal for congestion detection can result in delayed responses, as losses typically indicate a point of significant congestion. Packet loss detection, often based on timeouts, can be challenging to calibrate and may lead to unnecessary retransmissions. We categorize packet losses in two categories: losses because of severe congestion and losses due to networking failures. To differentiate the two types of losses and to improve reaction times, packet trimming, which only triggers for congestion drops, can be employed  (Adrian et al., 2022; Cheng et al., 2014; Handley et al., 2017; Olteanu et al., 2022). Trimming can be implemented on many existing switches (Adrian et al., 2022) and is starting to be supported by some switch vendors. It is also being pushed as a UE protocol feature (Consortium, [n. d.]). On current generation networks, REPS can be deployed with timeouts, and in next-generation augmented with trimming to distinguish between loss types (more details in Appendix A).
@@ 2.2. Load Balancing
is one of the most commonly used and simple load balancing mechanisms. It works by using a hashing function to randomly choose one of the available paths for a given packet (Hopps, 2009). The hashing functions usually takes as input 5 elements (aka five tuple) from the packet header: the protocol number, source address, destination address, source port, and destination port (some variations utilize only the four tuple without the protocol number). Recent approaches have proposed incorporating additional fields, such as the Time-to-Live (TTL) or the Flow Label (in IPv6), to further refine the hash calculation (Kabbani et al., 2014; Qureshi et al., 2022).
@@
is a value in the packet header that can be configured to be an input to the ECMP hashing function in switches. Such a value, which is set by the sender, allows it to alter a packet’s path in the network. Possible header fields to be used as an EV are the Source Port field in the packet header (Lu et al., 2018) or the Flow Label field in IPv6 (Qureshi et al., 2022). This is consistent with previous load balancing works and has already been demonstrated in practice (Kabbani et al., 2014; Qureshi et al., 2022; Vanini et al., 2017; Lu et al., 2018). We leverage EVs in REPS to improve load balancing and address the ECMP collision limitation, without needing to know the exact mapping between a packet’s EV and its resulting path.
@@
is a fixed-size set of EVs, since the number of possible values is constrained by the number of bits they can occupy in the packet header. For example, the source port field in a UDP header is assigned 16 bits, giving the EVS a size of 65536 possible values (excluding some reserved values) (Lu et al., 2018). While different numbers of bits could be allocated for the EVS, we analyze in Section 4.5.2 and Appendix B how many are required for optimal performance. It is generally advantageous for an algorithm to achieve good load balancing performance with a small EVS size since that often reduces the algorithm’s memory overhead. Note that the minimum effective EVS size also depends on the network topology: smaller topologies require fewer EVs to fully explore all available physical paths.
@@ 3.1. Core Logic: Path Exploration and Reuse
The circular buffer in REPS ensures that bursts of back-to-back ACKs with ”good” entropies are correctly cached and reused. Moreover, it guarantees stable load balancing in the case of failures as shown in Section 3.2. We use a circular buffer of 8 elements based on the bounds from Theorem 5.1.
@@ 3.2. Failure Mitigation: Freezing Mode
While this strategy could result in slightly worse load balancing (due to potentially reusing the same EV several times), it comes with the major benefit of guaranteeing that REPS will almost never pick the failing path again since the recent received EVs point to healthy paths. Considering the example above: by enabling freezing mode, the number of packets dropped decreases from over 120K packets to only about 1K.
To decide when to exit freezing mode, we either send probe packets to check the status of the failing paths or, if probing is not available, we simply exit freezing mode after a fixed amount of time. Once we exit freezing mode, we use random EVs to allow REPS to explore new paths and assess whether we detect new packets failing or not. This prevents REPS from getting stuck in a suboptimal state if the EVs in the buffer were all pointing to a failing path, a rare scenario that can theoretically happen with properly timed back-to-back network failures. Moreover, if REPS exits freezing mode before the issue is fully resolved, it will simply re-enter the mode shortly afterward with minimal impact on performance (as we will show in Section 4.3.3).
The intuition behind freezing mode is that once we suspect there is a failure, we want to start avoiding it as soon as possible. Interestingly, we observe that even if we enter freezing mode unnecessarily (i.e., by mistaking a congestion drop for a network failure), REPS would still load balance well, as discussed in Section 4.5.2 and Appendix A. This observation means that even if there is doubt about whether a real failure occurred, REPS can be conservative and can safely enter freezing mode.
@@
17:   $\mathit{exploreCounter}=\textit{NUM\_PKTS\_BDP}$
@@
6:   $\mathit{numberOfValidEVs}$--
@@
15:  if repsBuffer.isEmpty() or (numberOfValidEVs == 0 and not isFreezingMode) or exploreCounter then
16:   $dataPacket.ev=\textit{rand()}\,\%\,\mathit{EVS\_SIZE}$
17:   exploreCounter = $\max(\mathit{exploreCounter}-1,0)$
18:  else
19:   $dataPacket.ev=\textit{getNextEV()}$
20:  end if
21:end procedure
@@ 3.3. REPS Design Advantages
The simplicity of REPS enables cost-efficient hardware support, as it does not require any change of the packet headers format or existing network components. Moreover, its code is short and easy to implement and understand. REPS works best with per-packet ACKs, but we show in Section 4.5.1 that it still performs well even with ACK coalescing.
@@
Total (1 elements in buffer) 74 $\approx$10 bytes
@@
The general approach of REPS is that it only keeps track of good paths and avoids keeping statistics on congested or failing paths. This approach enables it to promptly load balance away from a congested link or failing link as, especially for the latter, it is never going to take a random guess once a link is failing. Any alternative method that tries to avoid selecting a failing path by tracking bad EVs would need to keep records of all the EVs that map to that path for a given flow, which would involve tracking not only the failing EVs but also all those still in flight.
@@ 4.1. Evaluation Setup
In large-scale simulations, we compare REPS with ECMP (Hopps, 2009), OPS (Dixit et al., 2013), PLB (Qureshi et al., 2022), MPRDMA (Lu et al., 2018), Flowlet Switching (Vanini et al., 2017), MPTCP (Raiciu et al., 2011), a bitmap approach where we keep per EV statistics similarly to STrack (Le et al., 2024), and adaptive RoCE by NVIDIA (NVIDIA, 2024). We note that this list includes both sender-based approaches and in-network switch-based approaches. We configure PLB to have more aggressive parameters similarly to FlowBender to improve its performance (Kabbani et al., 2014). For Flowlet Switching we set an aggressive flowlet timeout at half of the RTT. For MPTCP, we divide each message into 8 subflows and route each one individually similarly to using multiple QPs (Queue Pairs) (Gangidi et al., 2024).
@@
Regardless of the workloads that we discuss in Section 4.2, in the evaluation we simulate 3 different scenarios: (1) healthy symmetric network conditions, (2) asymmetric network conditions, which can be due to asymmetric path bandwidth or due to ECMP-hashed background traffic that increase load on specific paths (a scenario that can happen in incremental deployments of REPS), and (3) a network encountering various failures modes. We focus on the most relevant failure scenarios for real-world deployments (Gangidi et al., 2024) and we report the remaining ones in Appendix D.
@@
We implement REPS by extending the htsim packet-level network simulator (Handley et al., 2017). Our simulations consider different fat-tree topologies with 1024 nodes and 128 nodes and with different levels of oversubscription ranging from 1:1 (no oversubscription) to 4:1. We test 2- and 3-tier fat trees (TOR or Top-of-rack as T0, Aggregate as T1 and Core T2). Such topologies are commonly deployed in production datacenters designed for distributed training (Singh et al., 2015; Gangidi et al., 2024).
@@ 4.2. Workloads
consists of (1) incast, (2) permutation, and (3) tornado traffic patterns. Incast happens when multiple senders simultaneously send to one receiver. It is very common in storage workloads (Zhang et al., 2011; Chen et al., 2009) but also, with a small incast degree, in distributed training (Gangidi et al., 2024). In the permutation pattern, each node sends to a random receiver, and we ensure that each node is sending and receiving to exactly one node (Alizadeh et al., 2010). The tornado pattern is a special case of the permutation where each node sends to its ”twin” node in the other half of the tree. For example, with 128 nodes, node 0 would send to 64 and vice-versa, node 1 to 64 and so on. Tornado is an important worst case for load balancing, as each packet is required to traverse the full tree (Prisacari et al., 2013).
@@
We use real datacenter traces from similar previous work (Alizadeh et al., 2010; Zhang et al., 2017). We use a series of traces used for web search in production clusters. In such distribution the majority of flows are quite small (less than 100 KB) while a small number of flows are large. For each node we select randomly the receiver and run the simulation for 5 ms. More details are available in Appendix E.

### v6

@@ Abstract.
load balancing; packet spraying; out-of-order transport; RDMA; Ultra Ethernet
submissionid: 52ccs: Networks Data center networksccs: Networks Traffic engineering algorithmsjournalyear: 2026copyright: ccconference: 21st European Conference on Computer Systems; April 27–30, 2026; Edinburgh, Scotland Ukbooktitle: 21st European Conference on Computer Systems (EUROSYS ’26), April 27–30, 2026, Edinburgh, Scotland Ukdoi: 10.1145/3767295.3769320isbn: 979-8-4007-2212-7/26/04
@@ 1. Introduction
Modern AI training cluster networks draw on both high-performance computing (HPC) and cloud-native architectures (Hoefler et al., 2022b; Qian et al., 2024; Gangidi et al., 2024). These systems rely on Remote Direct Memory Access (RDMA) for low-latency, high-throughput communication. InfiniBand, a specialized RDMA transport, is widely used for its performance (36). More recently, operators have adopted commodity Ethernet fabrics, especially RoCEv2 (Association, 2024), which enable RDMA over standard hardware using off-the-shelf switches and NICs to reduce costs. However, as training clusters scale from around 10,000 endpoints (enough to train GPT-4 or Llama 3–scale models (Achiam et al., 2023; Dubey et al., 2024)) to over 100,000 nodes, both interconnects face fundamental limits in sustaining peak bandwidth due to:
@@
Thus, the community recognized the need for network stacks tailored to distributed training traffic while remaining compatible with commodity datacenter Ethernet infrastructure (Gangidi et al., 2024). Such proposals include SRD by Amazon (Shalev et al., 2020), Falcon by Google (Singhvi et al., 2025), TTPoE by Tesla (Motors, 2024), and Ultra Ethernet (UE) (Consortium, 2024), developed in collaboration between major tech players. Key open questions in these proposals are how to address load balancing and mitigate link failures.
@@
Our key insight is that the challenges of spraying-based load balancing can be addressed by adaptive packet spraying, paired with a transport layer that natively supports out-of-order packet delivery, as demonstrated in SRD (Shalev et al., 2020), UE (Lu et al., 2018; Hoefler et al., 2025), Falcon (Singhvi et al., 2025) and recent research work (Dixit et al., 2013; NVIDIA, 2024; Gangidi et al., 2024). Based on this insight, we design and contribute to the Ultra Ethernet Consortium (UEC) our decentralized load-balancing scheme, Recycled Entropy Packet Spraying aka REPS. REPS caches ”good” network paths in a circular buffer and quickly recovers (within a few RTTs) from network failures by adaptively discovering or freezing network paths. Figure 1 provides a schematic overview of REPS. The 1.0 Ultra Ethernet specification explicitly cites REPS as a reference load-balancing mechanism for Ultra Ethernet Transport (UET) (Ultra Ethernet Consortium, 2025).
Figure 1. Schematic of REPS. We do not show elements invalidation (for EV $5$ and 11) in the circular buffer after usage.
REPS does not require any special hardware support from switches beyond ECMP-based header hashing and ECN, which are standard features in modern switches (Hoefler et al., 2023; Nvidia, 2024). REPS requires only $\approx 25$ bytes of state per-connection regardless of the topology size, whereas MPTCP requires 368 extra bytes for 8 sub-flows (Lu et al., 2018).
We extensively evaluate REPS in simulation and augment the simulation findings with real hardware results in Section 4. We deploy REPS in a cluster with modified FPGA-based RDMA-capable NICs. In large-scale network simulations, REPS consistently outperforms state-of-the-art load balancing algorithms. REPS outperforms ECMP and OPS by up to $6\times$ and $1.25\times$ in symmetric networks, and by up to $5\times$ and $2\times$ in asymmetric networks, outperforming OPS by as much as $100\times$ during short-term transient link failures.
@@ 2. REPS Building Blocks
In this section, we introduce the fundamental building blocks necessary to understand REPS logic. We first describe key congestion signals, such as ECN (Floyd et al., 2001) and packet loss. We then introduce essential terms and concepts related to load balancing, including ECMP and EV. These concepts are largely established, but we revisit them here to provide clear terminology and a basis for the discussion that follows.
@@ 2.1. Congestion Signals
has long been a key indicator of severe congestion in networks (Floyd and Jacobson, 1993; Nichols and Jacobson, 2012). However, using packet losses as the sole signal for congestion detection can result in delayed responses, as losses typically indicate a point of significant congestion. Packet loss detection, often based on timeouts, can be challenging to calibrate and may lead to unnecessary retransmissions. We categorize packet losses in two categories: losses because of congestion and losses due to networking failures. To differentiate the two types of losses and to improve reaction times, packet trimming, which only triggers for congestion drops, can be employed (Adrian et al., 2022; Cheng et al., 2014; Handley et al., 2017; Olteanu et al., 2022; Li et al., 2025). Trimming can be implemented on many existing switches (Adrian et al., 2022) and is starting to be supported by some switch vendors. It is also being pushed as a UE protocol feature (Consortium, ). On current generation networks, REPS can be deployed with timeouts, and in next-generation augmented with trimming to distinguish between loss types (more details in Appendix A).
@@ 2.2. Load Balancing
is one of the most commonly used and simple load balancing mechanisms. It works by using a hashing function to randomly choose one of the available paths for a given packet (Hopps, 2009). The hashing function usually takes as input 5 elements (aka five tuple) from the packet header: the protocol number, source address, destination address, source port, and destination port (some variations utilize only the four tuple without the protocol number). Recent approaches have proposed incorporating additional fields, such as the Time-to-Live (TTL) or the Flow Label (in IPv6), to further refine the hash calculation (Kabbani et al., 2014; Qureshi et al., 2022).
@@
is a value in the packet header that can be configured to be an input to the ECMP hashing function in switches. This value, which is set by the sender, helps to determine a packet’s path through the network. Possible header fields to be used as an EV are the Source Port field in the packet header (Lu et al., 2018) or the Flow Label field in IPv6 (Qureshi et al., 2022). This is consistent with previous load balancing works and has already been demonstrated in practice (Kabbani et al., 2014; Qureshi et al., 2022; Vanini et al., 2017; Lu et al., 2018). REPS leverages entropy values (EVs) to improve load balancing and mitigate ECMP collision limitations, without requiring knowledge of the exact mapping from a packet’s EV to its resulting path. Switch hash functions can produce collisions: packets with distinct EVs, with all other header fields held constant (e.g., source and destination IPs), may still traverse the same physical path. The sender, including REPS, cannot directly infer whether changing the EV will change the path. Nevertheless, with a well-designed hash function, the induced path distribution should closely approximate a uniform random assignment, yielding near-optimal spreading.
@@
is a fixed-size set of EVs, since the number of possible values is constrained by the number of bits they can occupy in the packet header. For example, the source port field in a UDP header is assigned 16 bits, giving the EVS a size of 65536 possible values (excluding some reserved values) (Lu et al., 2018). While different numbers of bits could be allocated for the EVS, we analyze in Section 4.5.2 how many are required for optimal performance. It is generally advantageous for an algorithm to achieve good load balancing performance with a small EVS size since that often reduces the algorithm’s memory overhead. Note that the minimum effective EVS size also depends on the network topology: smaller topologies require fewer EVs to fully explore all available physical paths.
@@ 3.1. Core Logic: Path Exploration and Reuse
The circular buffer in REPS ensures that bursts of back-to-back ACKs with ”good” entropies are correctly cached and reused. Moreover, it guarantees stable load balancing in the case of failures as shown in Section 3.2. We use a circular buffer of 8 elements based on empirical evidence and the bounds derived from Theorem 5.1.
@@ 3.2. Failure Mitigation: Freezing Mode
While this strategy could result in slightly worse load balancing (due to potentially reusing the same EV several times), it comes with the major benefit of guaranteeing that REPS will almost never pick the failing path again since the recent received EVs point to healthy paths. Considering the example above: by enabling freezing mode, the number of drops decreases from over 120K packets to only about 1K.
To decide when to exit freezing mode, we either send probe packets to check the status of the failing paths or, if probing is not available, we simply exit freezing mode after a fixed amount of time. Once we exit freezing mode, we occasionally use random EVs to allow REPS to explore new paths and assess whether we detect new packets failing or not. This prevents REPS from getting stuck in a suboptimal state if the EVs in the buffer were all pointing to a failing path, a rare scenario that can theoretically happen with properly timed back-to-back network failures. Moreover, if REPS exits freezing mode before the issue is fully resolved, it will simply re-enter the mode shortly afterward with minimal impact on performance and a very small impact on extra packet drops (as we will show in Section 4.3.3).
The intuition behind freezing mode is that once we suspect there is a failure, we want to start avoiding it as soon as possible. Interestingly, we observe that even if we enter freezing mode unnecessarily (i.e., by mistaking a congestion drop for a network failure), REPS would still load balance well, as discussed in Section 4.5.2 and Appendix A. This observation means that even if there is doubt about whether a real failure occurred, REPS can be conservative and can safely enter freezing mode. Finally, we note that freezing mode is not crucial to the inner working of REPS, as it can work also without it. It merely adds more guarantees on top of it and improves performance in some failure cases. We showcase the impact of running REPS without freezing mode in Appendix C.4.
@@
17:   $\mathit{exploreCounter}=\textit{NUM\_PKTS\_CWND}$
@@
6:   $\mathit{numberOfValidEVs}=\mathit{numberOfValidEVs}-1$
@@
15:  if $\mathit{exploreCounter}>0$ then
16:   if $(\mathit{--exploreCounter}\%\mathit{REPS\_BUFFER\_SIZE})==0$ then
17:     $dataPacket.ev=\textit{rand()}\bmod\mathit{EVS\_SIZE}$
18:   end if
19:  else if repsBuffer.isEmpty() or (numberOfValidEVs == 0 and not isFreezingMode) then
20:   $\mathit{dataPacket.ev}$ = rand() % $\mathit{EVS\_SIZE}$
21:  else
22:   $\mathit{dataPacket.ev}$ = getNextEV()
23:  end if
24:end procedure
@@ 3.3. REPS Design Advantages
The simplicity of REPS enables cost-efficient hardware support, as it does not require any change of the packet headers format or existing network components. Moreover, its code is short and easy to implement and understand. REPS works best with per-packet ACKs, but we show in Section 4.5.1 that it still performs well even with ACK coalescing. While we focus on fat tree topologies in this paper, we believe REPS could work well even with different topologies (Kim et al., 2008; Hoefler et al., 2022a) with little to no adjustments required. Moreover, we expect REPS to work well even when using source-based routing algorithms where the NIC chooses directly the paths to use instead of relying on ECMP-hashing in switches. In this case, REPS would store the actual path ID for its EVs.
@@
Total (1 element in buffer) 74 $\approx$10 bytes
@@
The general approach of REPS is that it only keeps track of good paths and avoids keeping statistics on congested or failing paths. This approach enables it to promptly load balance away from a congested link or failing link as, especially for the latter, it is never going to take a random guess once a link is failing. Any alternative method that tries to avoid selecting a failing path by tracking bad EVs would need to keep records of all the EVs that map to that path for a given flow, which would involve tracking not only the failing EVs but also all those still in flight, resulting in excessive NIC’s memory usage.
@@ 4.1. Evaluation Setup
In large-scale simulations, we compare REPS with ECMP (Hopps, 2009), OPS (Dixit et al., 2013), PLB (Qureshi et al., 2022), MPRDMA (Lu et al., 2018), Flowlet Switching (Vanini et al., 2017), an MPTCP-like algorithm (Raiciu et al., 2011), a bitmap approach where we keep per EV statistics similarly to STrack (Le et al., 2024), and adaptive RoCE by NVIDIA (NVIDIA, 2024). We note that this list includes both sender-based approaches and in-network switch-based approaches. We configure PLB to have more aggressive parameters similarly to FlowBender to improve its performance (Kabbani et al., 2014). For Flowlet Switching we set an aggressive flowlet timeout at half of the RTT. For MPTCP, we divide each message into 8 subflows and route each one individually similarly to using multiple QPs (Queue Pairs) (Gangidi et al., 2024).
@@
Regardless of the workloads that we discuss in Section 4.2, in the evaluation we simulate 3 different scenarios: (1) healthy symmetric network conditions, (2) asymmetric network conditions, which can be due to asymmetric path bandwidth or due to ECMP-hashed background traffic that increase load on specific paths (a scenario that can happen in incremental deployments of REPS), and (3) a network encountering various failure modes. We focus on the most relevant failure scenarios for real-world deployments (Gangidi et al., 2024) and we report the remaining ones in Appendix C.
@@
We implement REPS by extending the htsim packet-level network simulator (Handley et al., 2017). Our simulations consider different fat-tree topologies with 1024 nodes and 128 nodes and with different levels of oversubscription ranging from 1:1 (no oversubscription) to 4:1. We test 2- and 3-tier fat trees (TOR or Top-of-rack as T0, Aggregate as T1 and Core T2). Such topologies are commonly deployed in production datacenters designed for distributed training (Singh et al., 2015; Gangidi et al., 2024). We use the larger topologies for the microbenchmark runs while the smaller topology for running the datacenter traces and AI collectives.
@@ 4.2. Workloads
consists of (1) incast, (2) permutation, and (3) tornado traffic patterns. Incast happens when multiple senders simultaneously send to one receiver. It is very common in storage workloads (Zhang et al., 2011; Chen et al., 2009) but also, with a small incast degree, in distributed training (Gangidi et al., 2024). In the permutation pattern, each node sends to a random receiver, and we ensure that each node is sending and receiving to exactly one node (Alizadeh et al., 2010). The tornado pattern is a special case of the permutation where each node sends to its ”twin” node in the other half of the tree. For example, with 128 nodes, node 0 would send to 64 and vice-versa, node 1 to 65 and so on. Tornado is an important worst case for load balancing, as each packet is required to traverse the full tree (Prisacari et al., 2013).
@@
We use real datacenter traces from similar previous work (Alizadeh et al., 2010; Zhang et al., 2017). We use a series of traces used for web search in production clusters. In such distribution the majority of flows are quite small (less than 100 KB) while a small number of flows are large. For each node we select randomly the receiver. More details are available in Appendix D.